wiki:UC20180507

Version 1 (modified by trac, 16 months ago) ( diff )

--

APO 3.5-m Users Committee Phonecon, 5/07/18

Attending: Nancy Chanover (NMSU), Mark Klaene (APO), Mary Beth Kaiser (JHU), John Bally (CU), Joanne Hughes Clark (Seattle U), Eric Bellm (UW), Kal Kadlec (UW), Sarah Tuttle (UW), Anne Verbiscer (UVa), Ben Williams (UW), Jon Holtzman (NMSU), Moire Prescott (NMSU), Russet McMillan (APO), Bill Ketzeback (APO), John Tobin (OU)


Identify minute-taker for this meeting: Nancy and Mark will do it - removing this item from the agenda in the future

Introduction of new members: Misty Bentz will be representing Georgia State going forward, although she was unable to attend today's call. Welcome Misty, and thank you Rachel for your service and participation on the Users Committee! Moire Prescott (NMSU) will be filling in for Jon Holtzman while he is on sabbatical. Welcome Moire!

User feedback and comments from institutional representatives:

  • NMSU - nothing to report
  • Colorado - There was a question from a user about why the summer shutdown will be so long this summer. See discussion below for more details.
  • Oklahoma - nothing to report
  • Wyoming - no report
  • Georgia State (report sent by email) - nothing to report
  • Washington - nothing to report
  • Seattle - no problems but Joanne spoke with Karen Kinemuchi re: issue with ARCTIC filters moving around in their holders - this is being addressed (see discussion of Telescope and Instrument report below)
  • UVa - Anne relayed a question from a user who is trying to reduce ARCTIC data and having pointing issues. The user observes a target, slews off target, and when they slew back, the pointing doesn’t match to within several arcseconds. John Bally reported that he has to redefine/update WCS in his images, e.g. by looking at 2MASS stars in the field and redefining WCS, basically give IRAF the coordinates of a reference star in both pixels and RA/Dec. This is a reduction fix, not a data acquisition fix. Russet said that the pointing should be accurate to within 4-5” but is not expected to be within 1 arcsec. One could use offsets to get stars back on same pixels if that’s important.
  • JHU - nothing to report
  • NAPG - no report

Telescope and Instruments Report:

Mark's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 4/1/18 through 5/1/18

0) Overview

Reasonably good weather this period but not great, lots more clouds than typical in April, we did have typical wind and dust though. DIS Blue Camera vacuum was serviced this past week to address ongoing scattered light issue.

1) Telescope

Telescope was operational with minimal issues. The time issues with TCC communications have been looked at and are being monitored. After some software updates the frequency may be slightly less but it did reoccur.

2) Instruments

DIS: Red camera scattered light is holding well , Blue deteriorated enough to cause us to service it again. DCam continues to give us issues but raising the absolute operating temperature seems to help some. Original DCam is still at Andor for servicing.

TripleSpec has been operational.

Agile has been operational this period.

Echelle is operational. A replacement slit viewer camera order from Fingerlakes has been ordered. It should be delivered in May so we can start software and hardware integration.

NIC-FPS is operational.

ARCTIC is operational. We will be working on improving filter wheel reliability over summer shutdown. Users should verify that their desired filter is in place.

3) ARCSAT

M1 and M2 have been reinstalled, collimation is in progress. Note this is the first time we have to re-collimate from scratch. We have some suggestions for DFM on improving their software to help with M2 focus issues and we hope to contract them to make these improvements.


Additional telescope and instrument discussion:

Weather wise it was a fairly typical April for us, no big surprises there. After the last DIS blue servicing we saw an increase in scattering as a result of what is thought to be contamination, hence we reserviced it again last week. Overall things have been going reasonably well. We saw one occurrence of the TCC timing issue; Fritz found a driver that needed to be updated and that has been done. This reduced the occurrence of these issues but we are not yet 100% convinced that it has gone away, so we will continue to monitor this issue.

The Dcam camera was repaired by Andor and it was delivered to APO today. The issues that we’ve been seeing recently have been with the spare, so if the repaired one looks good then we’ll send the spare back for repair. We did more engineering tests w/changing the Dcam temperature but that didn’t really improve things; the problem has to do with the current when it initially changes temperature - it is fine at first but then the readout noise continues to increase again.

The Fingerlakes camera that was purchased as a replacement for Ecam should show up this month but it will take some software development too. Hopefully it will be available and ready for use when we come out of summer shutdown.

The ARCTIC filter wheel movement repeatability issue is still being worked. Conor did some work on it but it's not quite where we want it to be yet. He will work on it more during summer shutdown. There are also some discussions underway among the engineering staff and obs-specs re: better filter inventory, tracking and management. This is quite involved currently and we are trying to make improvements, e.g. by making some new filter holders to reduce the number of filter swaps.

ARCSAT: M1 and M2 were reinstalled but Bill noticed that the coater failed to put a mark on secondary to indicate the center of the optical axis, which has ramifications for the collimation process. Kal and Bill have been discussing it and there is a plan going forward for the collimation process. Regarding the secondary focus issues, we will request software mods with DFM later this week. This may not address all the issues w/focussing but may make recovery easier and improve the reliability/repeatability of the focus mechanism.

Russet added that over the last few weeks the engineering staff has been working to understand a tracking oscillation in azimuth direction. They've been working with Fritz on the azimuth controller to reduce this. Currently it gets better in some parts of sky but worse in other parts. There have been a few mentions of this in the night logs, so users should report to the obs-specs if they notice anything unusual about tracking accuracy and we will continue to work on it. Some instruments lend themselves to this better than others - the effect will likely not be noticeable with DIS but will be detectable with ARCTIC.

Mark added that Fritz is retiring on July 1. They have been looking for a replacement for him and have had trouble finding appropriately qualified candidates. They recently reposted position at a mid-level and they hope to have someone new on board by July 1. Meanwhile they have started some conversations with Conor and John Brinkmann about how to cover any IT/computing/software needs if there is a gap between Fritz's retirement and the new hire.


Discussion of summer shutdown:

The summer shutdown is currently scheduled to run from July 5 to September 6, which is significantly longer than summer shutdowns have been in the recent past (although not very different from the summer shutdown durations of the early 2000's). The main reason for such a long shutdown is that we are doing work on both the secondary (M2) and tertiary (M3) mirrors, and that work can no longer be done on site (at Sunspot) and instead the mirrors must be sent to outside vendors. M2 and M3 are both made of Pyrex glass - they are early versions of Hextek mirrors. This Pyrex glass turns out to be very sensitive to acid stripping, which is what we have been using as a process for stripping and recoating these mirrors for the past 2 decades. We have now reached a point where the glass surface has been damaged, particularly on M3, and the way to correct that is to have the figure repolished, hence M3 will be repolished this summer to take out the roughness and reduce the scattered light from that mirror. That process will be done by Hextek, and then it will be sent to the coater. Both vendors have quoted conservative time estimates of 4 weeks, but the hope is that it will go faster than that. Other points mentioned in response to questions from the Users Committee were as follows:

  • M3 is flat so it does not make us nervous to have it repolished; it has been spec'ed to be returned with at least the same of flatness
  • M2 will just be stripped and recoated (not repolished) because it is not as degraded as M3 - it has not been recoated in 4 years due to problems we've had in the recent past with the chamber at Sunspot
  • The feeling is that this is important to do NOW because of the same degradation we saw on the ARCSAT mirror. In the last two summers Bill recoated the ARCSAT mirrors at Sunspot and within a week they could see a haze take over the coating and it got progressively worse over the year. Prior to sending out the ARCSAT mirror Bill talked to professional coaters and opticians and spoke w/Hextek; he learned that the 0.5m mirror and M3 are made of the same glass types and are reacting the same way. Although we don't have a quantitative measure of the degradation, with ARCSAT we did see a substantial reduction in throughput as a result of this haze.
  • Ultimately we will end up with M3 as a repolished mirror without the haze and with an enhanced Al coating w/SiO2 overcoat to protect it (better performance in blue, more durable). M2 will also have this enhanced Al coating.
  • We are not worrying about repolishing M2 this summer; that is another issue that will be tackled down the road. The primary is made of a different kind of glass and hence is not subject to this same degradation.

Nancy and Mark summarized by saying that they recognize that this long of a shutdown is painful, but it is also an investment in that we have not performed telescope improvements of this kind for many years, and hopefully will not have to do for quite a few years into the future after this summer. We scheduled the shutdown period to take place during the worst of the summer monsoon weather with the hope that this will not result in too many good nights that are sacrificed for these repairs/upgrades. Finally, the stripping, polishing (for M3) and recoating processes could go faster if the vendors are quicker than their most conservative estimates. If that happens we will notify users of a sooner-than-expected return to science ASAP.


Q2 3.5-m scheduling:

An update to the Q2 schedule was published earlier today. Russet notes that people who are interested in observing targets in the Kepler field will be thwarted by the lack of available time in Q3 (due to summer shutdown). Those people can request OPEN time in June, of which there is still some remaining, or they can propose for Q3 observations to be done in September.


Q2 ARCSAT scheduling:

We will not issue a call until we know when the telescope is back in service and available for science. We are planning for some preliminary shared-risk observing by people who had used ARCSAT fairly recently prior to its shutdown so that we can make a direct comparison of data quality before and and after the repolishing and recoating of the optics.


Q3 3.5-m scheduling:

Ben sent the Q3 allocations to all institutional schedulers; the requests are due on May 25. Institutional schedulers should keep a list of some programs that could be inserted into the schedule on relatively short notice if the shutdown ends earlier than anticipated.


Q3 ARCSAT scheduling:

There was some discussion of whether ARCSAT will be available during the 3.5m summer shutdown period. We are currently still working this out, but the answer is probably. It may be for a limited time, perhaps 3-4 weeks and maybe only on weeknights, because there won’t be day staff support or obs-specs who will be on site in the evenings, so they would need to provide support remotely. We will update the users once we have a better idea of the schedule.


NA2 Baffle Project

John Wilson provided an email report, which Nancy relayed to the Users Committee. The current status is that the UVa team has finished developing a Zemax model that matches the telescope SolidWorks model and they have set the secondary position in both models. Based on this optical model, when the NA2 baffle is ignored, quick analysis shows that the insides of the M3 baffle limits the FOV of the telescope to about 5.16 arcmin radial, which is good in that a redesigned NA2 baffle should therefore permit use of the full ARCTIC FOV. John will soon be working on the stray light analysis, including the vignetting effects due to the combined M2 and M3 baffle stack.


KOSMOS update

Kal and Sarah went to Tucson several weeks ago to check out the state of the instrument and work with the NOAO folks who are knowledgeable about KOSMOS. The good news is that KOSMOS also comes with a full test dewar and an engineering CCD, both of which can be immensely helpful to the UW team in getting ready to work on the actual instrument. The KOSMOS optics have been removed and were brought to Tucson to be cleaned. Kal is working with the NOAO folks to get the right number and size of crates - they are in the process of making sure they have everything inventoried while the optics get cleaned. Di Harmer will be working on the inventory this week so we hope things will move along relatively quickly in the short term.


TEWG Update

The Training and Education Working Group (TEWG) had its first telecon last week. The plan is now for the institutional reps on TEWG to provide their input regarding how they are currently using the 3.5m for teaching and training. There will also be a data-gathering phase where we solicit input from students regarding the effectiveness of their training sessions.


ACTION ITEMS:

Action items from previous meetings:

None.

New action items from this meeting:

  • Observers should report to the obs-specs if they notice anything unusual about the telescope tracking accuracy that may be caused by a tracking oscillation in azimuth direction. The effect will likely not be noticeable with DIS but will be detectable with ARCTIC, so not all observers may notice anything. But if you see something, say something!
  • Institutional schedulers should retain a list of programs that could be inserted into the schedule in late August/early September if the summer shutdown ends up being shorter than anticipated.

Next meeting:

The meeting schedule for the spring semester will be the first Monday of the month at 11:00 am Mountain Time. The next call will thus be June 4, 2018.

Future meetings are scheduled for:

  • July 9

We will evaluate everyone's availability (as well as the telecon schedule) once we get into summer months since we will need to contend with summer travel schedules as well as summer shutdown activities.

Final Comment:

This Wednesday (May 9) is the SDSS 20th anniversary of first light with the imager. To celebrate this milestone there will be a live Google Hangout, so anyone who was involved in SDSS during that time period may have received an announcement. Contact Mark for more information.

Note: See TracWiki for help on using the wiki.